OpenStudy (anonymous):

So I have g (x) = x squared + 2x Then I have g (a + h) - g (a) all over h, and I'm supposed to solve this I ended up with 2ah + 4a + 2h + 2a squared all over h, but the answer should be 2a + 2 + h. What did I do wrong?

OpenStudy (anonymous):

i believe this problem is finding the derivative of a function using the limit as h approaches 0. the answer should simply 2a + 2

OpenStudy (anonymous):

Well, that's not what my textbook says

OpenStudy (anonymous):

I just want to know how to get to the final answer, because at this point, I am completely stumped

OpenStudy (anonymous):

it would be: limit as h approaches 0 of: [ (a+h)^2 +2(a+h) - a^2 - 2a ] / h = limit as h approaches 0 of: [ a^2 +2ah + h^2 +2a + 2h - a^2 - 2a ] / h = limit as h approaches 0 of: [ 2ah + h^2 + 2h ] / h = limit as h approaches 0 of: 2a + h + 2 = 2a + 2

OpenStudy (anonymous):

my bad, just realized you aren't using limits. meaning it stops at the before last line, which is what your textbook says :)

OpenStudy (anonymous):

You are my hero right now...thank you for working this problem